# Neural Audio Codecs for Prompt-Driven Universal Source Separation

### v1

@@ 4.4 Results and Discussions
Table 4: Results: Generalization, Universal Source Separation (AudioCaps-test)
Model Separation
SI-SDR ($\uparrow$) ViSQOL ($\uparrow$)
AudioSep $\mathbf{-2.51^{\pm 12.14}}$ $\mathbf{2.44^{\pm 1.08}}$
AudioSep + dnr-v2 (zero-shot) $6.44^{\pm 11.48}$ $2.33^{\pm 1.08}$
CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.09^{\pm 11.62}}$ $2.24^{\pm 1.16}$
AudioSep + AudioCaps $9.17^{\pm 18.71}$ $2.29^{\pm 1.11}$
CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
Table 5: Results: Further Benchmarking on ESC-50, Clotho-v2, AudioSet-eval, VGGSound
Model Metric ($\uparrow$) Separation
ESC-50 Clotho-v2 AudioSet-eval VGGSound
AudioSep + dnr-v2 SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$
ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$
CodecSep + dnr-v2 SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$
ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$
Generalization across open-domain datasets.
Table 4 reports generalization results on the AudioCaps-test set derived from AudioSet. AudioSep benefits from distributional alignment, having been trained on diverse datasets, including AudioSet, and consequently achieves the strongest separation performance. CodecSep+dnr-v2 generalizes well to AudioCaps and outperforms AudioSep+dnr-v2 in SI-SDR while maintaining competitive ViSQOL scores. When retrained on AudioCaps, CodecSep again outperforms AudioSep in separation quality, demonstrating strong cross-domain robustness. Table 1 further supports this trend on the more challenging dnr-v2 test set, where CodecSep+AudioCaps outperforms AudioSep+AudioCaps in SI-SDR across all sources while maintaining comparable perceptual quality. However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
Further benchmarking on ESC-50, Clotho-v2, AudioSet-eval, and VGGSound (Table 5).
@@
Extending beyond AudioCaps and dnr-v2, we evaluate both systems on four additional open-domain benchmarks spanning environmental sound classification (ESC-50), audio captioning-style corpora (Clotho-v2), weakly labeled web-scale audio (AudioSet-eval), and visually grounded audio (VGGSound). Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all four datasets in both separation fidelity and perceptual quality. Taken together with the dnr-v2 analyses, these results indicate that CodecSep’s latent-domain masking—leveraging the codec’s structured, semantically aligned representations—confers robust cross-domain behavior while using $25\times$ less architecture-only compute than spectrogram separators. This combination of accuracy and efficiency is particularly attractive in practical deployments, where models must generalize to diverse sound taxonomies under strict compute and memory budgets.
@@
Robustness to prompt paraphrasing (Table 6) . To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}. This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference. We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts. Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
Table 6: Results: Impact of prompt ambiguity (dnr-v2-test)
Model Metric ($\uparrow$) Separation
Music Speech
Using Generic Prompts for Speech and Music
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$
@@
Using ambiguous prompts for Speech and Music
@@
Full inference complexity (Table 7) . We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep). Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep. Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space. Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting. Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth. Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
Relative–gain summaries (Appendix A).
For a compact view of percent improvements, Appendix A Table 8 reports relative gains of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text. In brief: on dnr-v2, CodecSep yields large improvements in SI-SDR with clear perceptual lifts; under paraphrased prompts, gains are smaller but remain positive, indicating robustness to lexical variation; across ESC-50, Clotho-v2, AudioSet-eval, and VGGSound, SI-SDR gains are consistent while ViSQOL deltas are modest; and when trained on AudioCaps, CodecSep maintains an advantage on AudioCaps-test and shows strong improvements when evaluated on dnr-v2.
Reconstruction study (Appendix B).
We probe source leakage with a single-source reconstruction diagnostic on dnr-v2-test : each input contains one target source and models are either prompted with the matching class/label (text-guided) or routed to the corresponding fixed head (non–text-guided). We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original. This is a leakage/consistency check—not a primary separation metric; full setup/results are in the Appendix B (Table 9).
@@ 5 Conclusion
Limitations and clarifications. (1) Codec/bandwidth. Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic. The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning. (2) Data and prompts. Training data scale and prompt diversity are modest relative to open-domain audio. As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains. (3) Temporal prompting. While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction. (4) Perceptual SFX quality. In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work. (5) Perceptual evaluation. Following CodecFormer, we report ViSQOL as a proxy for listening quality (predicting MOS-LQO on a 1–5 scale). Although we did not conduct a formal multi-listener study, we include real-world audio outputs for both CodecSep and AudioSep in the supplementary material, together with a Markdown summary of an informal author listening pass.
Future Works. We will scale CodecSep with higher-bandwidth/stereo codecs, broaden prompt coverage (including temporal and referring-expression cues), explore stronger self-alignment objectives, and investigate alternative conditioning (audio-only or multimodal) within the same FiLM-masking framework. Combined with its large compute reduction and code-stream compatibility, these extensions position CodecSep as a practical foundation for fast, universal, and deployable prompt-guided source separation.
Appendix A Discussion of Relative–Gain Summaries.

### v2

@@ 4.1 Results and Discussions
Further benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, & AudioCaps (cf. Table 4).
@@
Extending beyond dnr-v2, we evaluate both systems on five additional open-domain benchmarks spanning environmental sounds (ESC-50), audio captioning-style corpora (Clotho-v2, AudioCaps), weakly labeled web-scale audio (AudioSet), and visually grounded audio (VGGSound). Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all five datasets in both separation fidelity and perceptual quality.
Table 4: Results: Benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, AudioCaps
Model Metric ($\uparrow$) ESC-50 Clotho-v2 AudioSet VGGSound AudioCaps
AudioSep SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$ $6.44^{\pm 11.48}$
+ dnr-v2 ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$ $\mathbf{2.33^{\pm 1.08}}$
CodecSep SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$ $\mathbf{-6.09^{\pm 11.62}}$
+ dnr-v2 ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$ $2.24^{\pm 1.16}$
@@
Robustness to prompt paraphrasing (cf. Table 5) . To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}. This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference. We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts. Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
Full inference complexity (cf. Table 6) . We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep). Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep. Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space. Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting. Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth. Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
Additional experiments (cf. Appendix DF) .
For readability, three extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs. universal and paraphrased variants), training/evaluation splits, and metrics. (i) Generalization across open-domain datasets (cf. Appendix D): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL. When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality. The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR. (ii) Relative–gain summaries (cf. Appendix E): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf. Table 8). (iii) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided). We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original. This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix F (cf. Table 9).
@@
Table 5: Results: Using ambiguous prompts for Speech and Music (dnr-v2-test)
Model Metric ($\uparrow$) Music Speech
@@
Subjective evaluation (MOS–LQS).
We ran a human evaluation test with n=20n{=}20 participants on $20$ dnr-v2 $3$-stem test mixtures, comparing paired outputs from CodecSep+dnr-v2 and the official AudioSep model using fixed speech/music prompts and per-clip sfx prompts. Raters scored each stem independently in randomized order on the MOS–LQS scale (1=bad, 5=excellent); we report mean $\pm$ $1\sigma$. Overall, CodecSep scored $3.34^{\pm 1.00}$} vs. AudioSep $2.61^{\pm 1.04}$. By source, CodecSep achieved $3.17^{\pm 1.01}$} (music), $3.37^{\pm 0.97}$} (sfx), and $3.49^{\pm 1.00}$} (speech), while AudioSep obtained $2.49^{\pm 0.95}$, $2.84^{\pm 1.16}$, and $2.50^{\pm 1.02}$, respectively. These outcomes align with objective trends (SI-SDR/ViSQOL) and indicate consistent perceptual gains for CodecSep. Paired model outputs and reference stems are included in the supplementary materials for side-by-side listening.
Extensions: Multimodal prompting and high-bandwidth codecs.
Because conditioning enters only via a fixed-dimensional query embedding $e_{\tau}$ that drives FiLM in the masker, the architecture is agnostic to the prompt modality. Concretely, one can replace the text encoder with (i) an audio encoder to accept exemplar audio prompts $\big(e_{\tau}^{\text{aud}}\big)$, (ii) an image/vision–language encoder (e.g., CLIP-like) to accept image prompts $\big(e_{\tau}^{\text{vis}}\big)$, or (iii) a lightweight fusion (e.g., gated additive or attention pooling) of $\big(e_{\tau}^{\text{text}},e_{\tau}^{\text{aud}},e_{\tau}^{\text{vis}}\big)$ to support mixed prompts—all without modifying the masker or the codec. Likewise, the separation stack is codec-agnostic: upgrading the frozen backbone from 16 kHz to full-band 44.1/48 kHz simply changes the latent shape $(d,T)$, primarily along the feature axis rather than the time dimension; the FiLM-conditioned masker and training objective remain unchanged.
@@ 5 Conclusion
Limitations and clarifications. (1) Codec/bandwidth. Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic. The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning. (2) Data and prompts. Training data scale and prompt diversity are modest relative to open-domain audio. As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains. (3) Temporal prompting. While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction. (4) Perceptual SFX quality. In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work.
Future Work. We will (i) scale the frozen codec backbone to higher bandwidths and stereo (44.1/48 kHz), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
Appendix A Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
Problem setup and pipeline contrast.
Let $x(t)\in\mathbb{R}$ be a mono mixture with sources $\{y_{s}(t)|s\in\mathcal{S}\}$, where $x(t)$ is expressed as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$.
Spectrogram-domain (AudioSep):
$x\;\xrightarrow{\text{STFT}}\;X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\odot|X|\exp(\angle X+\angle\hat{M}_{s})\;\xrightarrow{\text{ISTFT}}\;\tilde{y}_{s}(t),$
where $g(\cdot,e_{\tau})$ denotes a FiLM-conditioned, complex-domain spectrogram separator that predicts a magnitude mask $|\hat{M}_{s}|\in[0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$, conditioned jointly on the mixture spectrogram $X$ (obtained by the STFT of $x$) and the text embedding $e_{\tau}$. The predicted magnitude mask and phase residual are then applied element-wise to form the source spectrogram $\tilde{Y}_{s}$ which is subsequently transformed back to the time domain via ISTFT to obtain $\tilde{y}_{s}(t)$.
CodecSep (NAC latent-domain):
Dimensionality reduction and compression.
Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors. For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$. A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention. Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs.
A.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
$Enc(.):x\mapsto Z\in\mathbb{R}^{d\times T}.$
$\tilde{Z}_{k}\;=\;M_{k}\odot Z,\qquad\tilde{y}_{k}\;=\;Dec(\tilde{Z}_{k}),$
$Mask(.,.):\;Z\in\mathcal{M}_{\text{latent}}\;\to\;\hat{Z}_{k},$
$g(.,.):\;X\in\mathbb{C}^{F\times T}\;\to\;\hat{Y}_{k},$
Hierarchical (RVQ) structure that benefits separation. The codec applies Residual Vector Quantization (RVQ) to $Z$, producing discrete codes $A=[a_{t}\in[K]^{N_{q}}]_{t=1}^{T}$ with $K=1024$ and $N_{q}$ quantizers. Codebook lookup yields embeddings
The overall loss is
A.2 Transformer-based masker vs. decoder-style generation (specific design choice).
  • Direct leverage of codec-disentangled latents. DAC/RVQ pretraining organizes $Z$ into a coarse-to-fine, semantically disentangled manifold (periodicity, timbre, transients). The masker exploits this structure by gating along these axes to isolate sources, rather than using $Z$ merely as input to a decoder that generates new latents/waveforms. This turns separation into selection on a well-structured space, improving identifiability while reducing parameters and data demands.
A.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
We deliberately integrate Feature-wise Linear Modulation (FiLM) Perez et al. [2018] inside the transformer masker to inject text semantics while preserving the codec manifold and keeping $Enc(\cdot)$$Dec(\cdot)$ frozen.
Targeted placement (masker, mid-layers).
Lightweight computation (overhead and parameterization).
Non-iterative inference (single forward pass).
Manifold preservation and stability.
A.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
Why we evaluate on continuous latents $Z$.
We choose to perform—and therefore report—separation on the continuous DAC encoder latents $Z=Enc(x)\in\mathbb{R}^{d\times T}$ for the following concrete reasons:
  1. 1.
  2. 2.
  3. 3.
How to extend the same model to discrete code streams.
$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big(a_{t}^{(i)}\big),$ (3)
$E$ $=[e_{t}]_{t=1}^{T}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (4)
$\tilde{E}_{s}\;=\;M_{s}\odot E\;\approx\;M_{s}\odot Z\;=\;Z_{s}$
Embedding alignment (CodecFormer-EL Yip et al. [2024a]). To tighten $\tilde{E}_{s}\approx Z_{s}$, we can optimize an embedding-level consistency loss as in CodecFormer-EL Yip et al. [2024a]instead of our SI-SDR objective:
$\mathcal{L}_{\mathrm{emb}}\;=\;\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$
A.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
Pipeline comparison (server-side separation).
Traditional (spectrogram) pipeline:
CodecSep pipeline:
Complexity accounting.
In the common (code-stream) case, CodecSep removes both decode and re-encode on the server. Moreover, within the separator, CodecSep operates on $Z/E$ with $|Z|\ll|X|$ (cf. Sec. A), reducing activation memory and bandwidth throughout attention and MLP blocks.
Interface compatibility with codec bitstreams.
Why spectrogram-domain systems incur extra overhead.
Operational advantages of CodecSep.
  • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec. A.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
Appendix B Dataset Details
B.1 Divide and Remaster v 2.0 (dnr-v2)
dnr-v2 Petermann et al. [2022] dataset consists of $60s$-duration artificial mixtures of speech, music, and SFX sampled from LibriSpeech Panayotov et al. [2015], Free Music Archive (FMA) Defferrard et al. [2017], and Freesound Dataset $50K$ (FSD50K50K) Fonseca et al. [2022], respectively. It includes $3,406$ ($56.7hrs$) training, $487$ ($8.13hrs$) validation, and $973$ ($16.22hrs$) test mixtures, each provided with its three individual source audios. The mixtures are generated by normalizing each source to fixed Loudness Units Full-Scale (LUFS) levels: 17-17 dB (speech), 24-24 dB (music), and 21-21 dB (SFX), with $\pm 2$ dB random perturbations. Any source exceeding a peak threshold is normalized to $0.5$ dB. The sources are mixed and normalized to 27-27 dB LUFS with additional random perturbations. The validation and test sets are trimmed for silence and split into $5s$ or $10s$ segments. Segments where sources are present for less than $50\%$ of the duration are removed, resulting in $2,852$ ($\approx 3.96hrs$) validation and $1,840$ ($\approx 5.11hrs$) test mixtures.
B.2 Open-Domain Benchmarks
We benchmark on five open-domain datasets spanning captioned audio, environmental sounds, and large multi-event corpora: AudioCaps Kim et al. [2019], an AudioSet-derived collection of >46>46k $10$ s YouTube clips paired with human-written captions describing the dominant sound events (used by us to synthesize training and test mixtures); ESC-50 Piczak , a curated environmental sound dataset of 2,0002{,}000 clips ($5$ s each) organized into $50$ classes with $40$ examples per class across five meta-categories (animals, natural, human non-speech, domestic, exterior/urban); Clotho-v2 Drossos et al. [2020], 6,9746{,}974 audio samples (15153030 s) each annotated with five human captions (8–20 words) covering open-domain events; AudioSet Gemmeke et al. [2017], the evaluation split of AudioSet comprising human-labeled $10$ s YouTube clips over an ontology of $632$ audio event classes in a multi-label setting; and VGGSound Chen et al. [2020a], an AudioSet-derived audio–visual corpus with 550+550{+} hours of $10$ s segments covering a wide variety of everyday sound categories. For AudioCaps we form both training and testing mixtures (same scale of test data as dnr-v2) by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
Appendix C Training Details
The complete model, including the query module $query(.)$, is trained for $400K$ iterations with DAC Kumar et al. [2023] and CLAP Wu et al. [2023] modules frozen. Validation is conducted every $5K$ iterations and test every $10K$ iterations. We use ADAM Kingma and Ba [2017] as our optimizer and train with a batch size of $4$ examples, each $2$ seconds in duration, and a learning rate of $1.5e^{-4}$ on a single 2424GB NVIDIA A-$30$ GPU. Training employs a $ReduceLRonPlateau$ Mukherjee et al. [2019] scheduler, which reduces the learning rate by a factor of $0.5$ if the validation loss does not improve for two consecutive validation checks. We train two versions of CodecSep, one using the dnr-v2 dataset and the other using AudioCaps, to evaluate performance across different training distributions. We refer to these models using the suffixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
Appendix D Cross-Benchmark Performance: AudioCaps and dnr-v2.
Table 7(a) reports generalization results on the AudioCaps-test set derived from AudioSet. AudioSep benefits from distributional alignment, having been trained on diverse datasets, including AudioSet, and consequently achieves the strongest separation performance. CodecSep+dnr-v2 generalizes well to AudioCaps and outperforms AudioSep+dnr-v2 in SI-SDR while maintaining competitive ViSQOL scores. When retrained on AudioCaps, CodecSep again outperforms AudioSep in separation quality, demonstrating strong cross-domain robustness. Table 7(b) further supports this trend on the more challenging dnr-v2 test set, where CodecSep+AudioCaps outperforms AudioSep+AudioCaps in SI-SDR across all sources while maintaining comparable perceptual quality. However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
Table 7: Generalization and transfer results for universal source separation.
(a) Generalization on AudioCaps-test.
Model Separation
SI-SDR ($\uparrow$) ViSQOL ($\uparrow$)
AudioSep $\mathbf{-2.51^{\pm 12.14}}$ $\mathbf{2.44^{\pm 1.08}}$
AudioSep + dnr-v2 (zero-shot) $6.44^{\pm 11.48}$ $2.33^{\pm 1.08}$
CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.09^{\pm 11.62}}$ $2.24^{\pm 1.16}$
AudioSep + AudioCaps $9.17^{\pm 18.71}$ $2.29^{\pm 1.11}$
CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
(b) Transfer to dnr-v2-test when trained on AudioCaps (zero-shot on dnr-v2).
Model Metric ($\uparrow$) Music Speech Sfx
(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
Appendix E Discussion of Relative–Gain Summaries.